# Information Re-Organization Improves Reasoning in Large Language Models

### v1

@@ Abstract
Improving the reasoning capabilities of large language models (LLMs) has attracted considerable interest. Recent approaches primarily focus on improving the reasoning process to yield a more precise final answer. However, in scenarios involving contextually aware reasoning, these methods neglect the importance of first identifying logical relationships from the context before proceeding with the reasoning. This oversight could lead to a superficial understanding and interaction with the context, potentially undermining the quality and reliability of the reasoning outcomes. In this paper, we propose an information re-organization (InfoRE) method before proceeding with the reasoning to enhance the reasoning ability of LLMs. We first perform a re-organization processing of the contextual content, e.g., documents or paragraphs, to obtain logical relationships. Then, we utilize the re-organized information in the reasoning process. This enables LLMs to deeply understand the contextual content by clearly perceiving these logical relationships. To demonstrate the effectiveness of our approach in improving the reasoning ability, we conduct experiments using Llama2-70B, GPT-3.5, and GPT-4 on various contextually aware multi-hop reasoning tasks. Using only a zero-shot setting, our method achieves an average improvement of 3% across all tasks, highlighting its potential to improve the reasoning performance of LLMs. Our source code is available at https://github.com/hustcxx/InfoRE.
@@ 1 Introduction
Large language models (LLMs) demonstrate powerful generative capabilities and achieve remarkable performance across a range of linguistic tasks Brown et al. (2020); Touvron and et al. (2023); OpenAI (2023). However, their capabilities in performing complex reasoning tasks — an essential aspect of advanced language understanding and intelligent decision-making — still present substantial challenges Arkoudas (2023); Blair-Stanek et al. (2023). This has spurred researchers to explore innovative strategies Wei et al. (2022); Yao et al. (2023); Besta et al. (2023) to improve the reasoning capabilities of these models.
Figure 1: InfoRE (Ours) vs existing methods for improving the reasoning of LLMs. In contrast to the existing methods that primarily focus on the reasoning process, our InfoRE emphasizes the re-organization of context information. The [TEXT] in italics indicates that it is optional in the reasoning process.
@@
Recently, diverse methods have been developed to enhance the reasoning ability of LLMs. For example, a notable method Chain-of-Thought (CoT) Wei et al. (2022), incorporates a series of intermediate reasoning steps into the reasoning. CoT Wei et al. (2022) allows for a more transparent and understandable path to the final answer, making it easier to follow the logic behind the conclusion. Building upon this foundation, subsequent approaches such as Tree of Thoughts (ToT) Besta et al. (2023) and Graph of Thoughts (GoT) Besta et al. (2023) are proposed to further refine the reasoning steps and enhance the accuracy and reliability of LLMs. Different from the sequential intermediate steps of Chain-of-Thought Wei et al. (2022), Tree of Thoughts Besta et al. (2023) and Graph of Thoughts Besta et al. (2023) model the problem solving process into structured tools of tree and graph, respectively. A critical observation is that these existing approaches primarily focus on improving the reasoning process of large language models, as shown in Figure 1 (left). However, in scenarios involving contextually aware reasoning, it is equally important to first identify logical relationships of context before proceeding with the reasoning, not just improve their reasoning process. This is because logical relationships, such as parallelism, causal connections, contrasts, etc., are essential elements of reasoning Braine (1978). Nevertheless, these existing methods often neglect this crucial step. Such an oversight can lead to a superficial understanding and interaction with the context, potentially undermining the quality and reliability of the reasoning results.
@@
Inspired by the fact that when faced with context-aware reasoning tasks humans often first re-organize existing contextual information to uncover the logical relationships and enhance their understanding of the context, we propose an information re-organization (InfoRE) method, to ground reasoning by the re-organized information. As shown in Figure 1 (right), different from previous methodologies that primarily focus on refining the reasoning steps to enhance the reasoning capabilities of LLMs, our approach takes a novel direction of context re-organization. We emphasize the utilization of re-organized contextual content to explicitly present the logical relationships that are often implicit within the plain text, promoting more effective reasoning. Specifically, we first perform a re-organization processing of the contextual content into a MindMap structure Buzan et al. (2010), uncovering logical relationships implicit within the context. We employ this structure for re-organized information because it is rich in logical relationships and encompasses multi-hop connections. Then, we utilize the re-organized context to reason. This enables LLMs to deeply understand the contextual content by clearly perceiving these logical relationships, facilitating the quality and reliability of reasoning. Besides, our information re-organization method can be integrated with existing prompt methods, like CoT Wei et al. (2022), to further improve the reasoning ability of LLMs. To verify the efficacy of our proposed InfoRE method, we conduct experiments using Llama2-70B Touvron and et al. (2023), GPT-3.5 Brown et al. (2020), and GPT-4 OpenAI (2023) on various contextually aware multi-hop reasoning tasks, including claim verification Guo et al. (2021), question answering Perez et al. (2020), and reading comprehension Dua et al. (2019). Using only a zero-shot setting, our method achieves an average improvement of 3% across all tasks, highlighting its potential to improve the reasoning performance of LLMs.
@@
  • In contrast to existing methodologies that primarily focus on refining the reasoning steps to enhance the reasoning capabilities of LLMs, we take a novel direction of context re-organization.
  • We first perform a re-organization processing of the contextual content to uncover the logical relationships that encompass multi-hop connections of the existing context, then to facilitate the reasoning of LLMs.
@@ 2 Related Work
Large language models (LLMs) Touvron and et al. (2023); Brown et al. (2020); OpenAI (2023), have revolutionized the field of natural language processing (NLP) and demonstrate remarkable proficiency across a range of linguistic tasks. To further improve the reasoning ability of LLMs has attracted considerable interest. In-context learning Brown et al. (2020), as a promising approach to enhance the reasoning abilities of LLMs, has been verified in mathematical reasoning task Wei et al. (2022). In addition, Chain of Thought (CoT) Wei et al. (2022) incorporates a coherent series of intermediate reasoning steps to improve the reasoning ability of LLMs. Following this paradigm, Tree-of-Thoughts (ToT) Yao et al. (2023) and Graph-of-Thoughts (GoT) Besta et al. (2023) have also been proposed to focus on improving the structure of the intermediate chain. Then self-consistency Wang et al. (2022) and plan-to-solve Wang et al. (2022) focus on the reliability of the chain. Recently, the step-back prompting Zheng et al. (2023) is proposed, which obtains the high-level concept and first principles from instances by abstraction in the first step, then guides the reasoning of LLMs with the obtained concept and principles. For tasks that require multi-hop reasoning, current approaches Pan et al. (2023); Radhakrishnan et al. (2023) generally decompose multi-hop problems into simpler sub-tasks problems relying on the in-context learning ability of LLMs. In contrast to existing methods, we take a novel direction of context re-organization to enhance the reasoning capabilities of LLMs for contextually aware reasoning tasks.
@@
Information re-organization is a technique that leverages other structures to improve the clarity and comprehensibility of information. Moreover, information re-organization can also reveal conceptual relationships implicit in the original textual text. The graph, as the most common re-organization structure, is often used to enhance contextual representation in a variety of tasks. For example, BASS Wu et al. (2021) re-organizes the original document to a semantic graph to improve the summary generation. DGM Ouyang et al. (2021) re-organizes the context to an explicit discourse graph in the reading comprehension task. MindMap Buzan et al. (2010) is a powerful structure method for representing knowledge and concepts, it can be used to construct a hierarchical abstraction of natural language text Elhoseiny and Elgammal (2014). Previous method Dalamagas et al. (2010) uses it to organize a large amount of scientific material to enhance the clarity of the text material. In this paper, we use it as the structure of re-organized information to uncover the logical relationships and multi-hop connections implicit within the plain context. Then we use this re-organized information to further improve the reasoning ability of LLMs on multi-hop reasoning tasks.
@@ 3 Methodology
In this section, we first give a formulation of a context-aware reasoning task in §3.1 and then describe our method in detail. As shown in Figure 2, our method consists of two steps, an information re-organization step §3.2 and a reasoning step with the re-organized context information §3.3.

### v2

@@ Abstract
Improving the reasoning capabilities of large language models (LLMs) has attracted considerable interest. Recent approaches primarily focus on improving the reasoning process to yield a more precise final answer. However, in scenarios involving contextually aware reasoning, these methods neglect the importance of first identifying logical relationships from the context before proceeding with the reasoning. This oversight could lead to a superficial understanding and interaction with the context, potentially undermining the quality and reliability of the reasoning outcomes. In this paper, we propose an information re-organization (InfoRE) method before proceeding with the reasoning to enhance the reasoning ability of LLMs. Our re-organization method involves initially extracting logical relationships from the contextual content, such as documents or paragraphs, and subsequently pruning redundant content to minimize noise. Then, we utilize the re-organized information in the reasoning process. This enables LLMs to deeply understand the contextual content by clearly perceiving these logical relationships, while also ensuring high-quality responses by eliminating potential noise. To demonstrate the effectiveness of our approach in improving the reasoning ability, we conduct experiments using Llama2-70B, GPT-3.5, and GPT-4 on various contextually aware multi-hop reasoning tasks. Using only a zero-shot setting, our method achieves an average absolute improvement of 4% across all tasks, highlighting its potential to improve the reasoning performance of LLMs ***https://github.com/hustcxx/InfoRE.
@@ 1 Introduction
Large language models (LLMs) demonstrate powerful generative capabilities and achieve remarkable performance across a range of linguistic tasks [1, 2, 3]. However, their capabilities in performing complex reasoning tasks — an essential aspect of advanced language understanding and intelligent decision-making — still present substantial challenges [4, 5]. This has spurred researchers to explore innovative strategies [6, 7, 8] to improve the reasoning capabilities of these models.
@@
Recently, diverse methods have been developed to enhance the reasoning ability of LLMs. For example, a notable method Chain-of-Thought (CoT) [6], incorporates a series of intermediate reasoning steps into the reasoning. CoT [6] allows for a more transparent and understandable path to the final answer, making it easier to follow the logic behind the conclusion. Building upon this foundation, subsequent approaches such as Tree of Thoughts (ToT) [8] and Graph of Thoughts (GoT) [8] are proposed to further refine the reasoning steps and enhance the accuracy and reliability of LLMs. Different from the sequential intermediate steps of Chain-of-Thought [6], Tree of Thoughts [8] and Graph of Thoughts [8] model the problem solving process into structured tools of tree and graph, respectively. A critical observation is that these existing approaches primarily focus on improving the reasoning process of large language models, as shown in Figure 1 (left). However, in scenarios involving contextually aware reasoning, it is equally important to first identify logical relationships of context before proceeding with the reasoning, not just improve their reasoning process. This is because logical relationships, such as parallelism, causal connections, contrasts, etc., are essential elements of reasoning [9]. Nevertheless, these existing methods often neglect this crucial step. Such an oversight can lead to a superficial understanding and interaction with the context, potentially undermining the quality and reliability of the reasoning results.
Figure 1: InfoRE (Ours) vs existing methods. In contrast to the existing methods that primarily focus on the reasoning process, our InfoRE emphasizes the re-organization of context information.
@@
Inspired by the fact that when faced with context-aware reasoning tasks humans often first re-organize existing contextual information to uncover the logical relationships, eliminate noises, and enhance their understanding of the context, we propose an information re-organization (InfoRE) method, to ground reasoning by the re-organized information. As shown in Figure 1 (right), different from previous methodologies that primarily focus on refining the reasoning steps to enhance the reasoning capabilities of LLMs, our approach takes a novel direction of context re-organization. We emphasize the utilization of re-organized contextual content to explicitly present the logical relationships that are often implicit within the plain text, promoting more effective reasoning. Specifically, our re-organization method comprises two operations: extraction and pruning. The extraction first uncovers the implicit logical relationships within the contextual content by transforming the content into a MindMap structure [10]. We employ this structure because it is rich in logical relationships and encompasses multi-hop connections. Pruning is then used to further minimize noise that is irrelevant to the reasoning objective. The pruning operation uses a pre-trained BERT [11] based model trained with reinforcement learning (RL). Finally, we utilize the re-organized context to reason. This enables LLMs to deeply understand the context by clearly perceiving these logical relationships, facilitating the quality and reliability of reasoning. Besides, our information re-organization method can be integrated with existing prompt methods, like CoT [6], to further improve the reasoning ability of LLMs. To verify the efficacy of our proposed InfoRE method, we conduct experiments using Llama2-70B [2], GPT-3.5 [1], and GPT-4 [3] on various contextually aware multi-hop reasoning tasks, including claim verification [12], question answering [13], and reading comprehension [14]. Using only a zero-shot setting, our method achieves an average improvement of 4% across all tasks, highlighting its potential to improve the reasoning performance of LLMs.
@@
  • In contrast to existing methods that primarily focus on refining the reasoning steps to enhance the reasoning capabilities of LLMs, we take a novel direction of context re-organization.
  • The re-organization method initially uncovers the logical relationships that encompass multi-hop connections in the contextual content by extraction, and subsequently minimizes noise by pruning.
@@ 2 Related Work
Large language models (LLMs) [2, 1, 3], have revolutionized the field of natural language processing (NLP) and demonstrate remarkable proficiency across a range of linguistic tasks. To further improve the reasoning ability of LLMs has attracted considerable interest. In-context learning [1], as a promising approach to enhance the reasoning abilities of LLMs, has been verified in mathematical reasoning task [6]. In addition, Chain of Thought (CoT) [6] incorporates a coherent series of intermediate reasoning steps to improve the reasoning ability of LLMs. Following this paradigm, Tree-of-Thoughts (ToT) [7] and Graph-of-Thoughts (GoT) [8] have also been proposed to focus on improving the structure of the intermediate chain. Then self-consistency [15] and plan-to-solve [15] focus on the reliability of the chain. Recently, the step-back prompting [16] is proposed, which obtains the high-level concept and first principles from instances by abstraction in the first step, then guides the reasoning of LLMs with the obtained concept and principles. For tasks that require multi-hop reasoning, current approaches [17, 18] generally decompose multi-hop problems into simpler sub-tasks problems relying on the in-context learning ability of LLMs. In contrast to existing methods, we take a novel direction of context re-organization to enhance the reasoning capabilities of LLMs for contextually aware reasoning tasks.
@@
Information re-organization is a technique that leverages other structures to improve the clarity and comprehensibility of information. Moreover, information re-organization can also reveal conceptual relationships implicit in the original textual text. The graph, as the most common re-organization structure, is often used to enhance contextual representation in a variety of tasks. For example, BASS [19] re-organizes the original document to a semantic graph to improve the summary generation. DGM [20] re-organizes the context to an explicit discourse graph in the reading comprehension task. MindMap [10] is a powerful structure method for representing knowledge and concepts, it can be used to construct a hierarchical abstraction of natural language text [21]. Previous method [22] uses it to organize a large amount of scientific material to enhance the clarity of the text material. In this paper, we use it as the structure of re-organized information to uncover the logical relationships and multi-hop connections implicit within the plain context. Then we use this re-organized information to further improve the reasoning ability of LLMs.
@@ 3 Methodology
In this section, we first give a formulation of a context-aware reasoning task in §3.1 and then describe our method in detail. As shown in Figure 2, the framework of our method consists of two components, an information re-organization §3.2 and a reasoning step using the re-organized context §3.3.
Figure 2: Illustration of our information re-organization method with two modules: 1)Information Re-Organization, which includes logic relationship extraction and noise pruning. 2) Reasoning using re-organized context. The re-organized context in black italicized text is relevant to the question.
3.1 Task Formulation
Given a sample $(c,q,a)$ from sample space $(\mathcal{C},\mathcal{Q},\mathcal{A})$, where $c$ denotes gold context, $q$ denotes question, the task aims to obtain answer $a$ to the question $q$, based on gold context $c$ with a large language model. In particular, it requires multi-hop reasoning to get the answer $a$ to question $q$.
3.2 Information Re-Organization
The purpose of the information re-organization is to obtain the logical relationships from the context and minimize the noise irrelevant to the question. This goal is achieved through two operations: extraction and pruning.
3.2.1 Extraction
For a question $q$, if the derivation of its answer $a$ relies on context $c$, then a deeper understanding of $c$ is crucial. Logical relationships, such as parallelism, causal connections, contrasts, etc., are essential elements of understanding and reasoning [9]. However, the logical relationships in the plain context are often implicit. Therefore, we perform a extraction operation on the plain context to uncover the logical relations in it using a language model. The process can be defined as:
$g=f_{\theta}(c,q,P_{in})$ (1)
where $f_{\theta}(\cdot)$ is a language model parameterized by $\theta$, $P_{in}$ is input task prompt. The specific task prompt $P_{in}$ used in our paper is displayed in the Appendix A. We perform Equation 1 for each $c$ in the sample space to obtain the extracted context space $\mathcal{G}$. We use the MindMap [10] structure to display the reorganized content, because it serves as a powerful structure for representing knowledge, concepts, and perspectives, contains not only logical relationships but also multi-hop connections.
@@ 3.2.1 Extraction
3.2.2 Pruning
As described in Section 3.2, the extracted context $g\in\mathcal{G}$ contains various logical relationships and rich attributes. However, not all logical relationships and attributes help answer the question $q$. On the contrary, some may even interfere with the response to the question. For example, consider the question in Figure 2, "Where did the producer of Julius Caesar study or work?". In this case, the content "Julius Caesar $\rightarrow$ Production Company" is a distracting element, and "Julius Caesar $\rightarrow$ Adaptation" is irrelevant to the question.
Figure 3: Illustration of Pruning model. The representation of [CLS] is used to obtain action probabilities.
To further reduce the interference of distracting or irrelevant logical relations and attributes on the retrieval of answers for question $q$, we use a pruning model trained through reinforcement learning (RL). The pruning model is based on the pre-trained BERT [11], as shown in Figure 3. Its input consists of concatenated logical relationships, their corresponding attribute values from $g$, and the question $q$. For example, to prune the relation Julius Caesar $\rightarrow$ Adaptation, the input is: [CLS] Julius Caesar Adaptation Play by Shakespeare [SEP] Question [SEP]. We give a detailed demonstration of input format in Appendix B.
RL Formulation
We formulate the pruning policy model optimization as an RL problem and employ proximal policy optimization (PPO) [23]. The action is keeping or deleting the logical relations. The policy decides the action probability given the question and $g$. We fine-tune the policy model $\pi$ by optimizing the reward $r$:
$\mathbb{E}_{\pi}[r]=\mathbb{E}_{\boldsymbol{g}\sim\mathcal{G},\boldsymbol{q}\sim\mathcal{Q},\boldsymbol{z}\sim\pi(\cdot\mid\boldsymbol{x},\boldsymbol{q})}[r(\boldsymbol{z},\boldsymbol{q})]$ (2)
Reward Function
Our goal is to maximize LLM’s generation toward the desired target by an alignment measure $\mathcal{R}$, and we use this as a reward. In this paper, the alignment metric we have chosen is the F1 score. To keep the policy network $\pi$ from moving too far from the old result, we add a clipped term in the reward. Therefore, the final reward becomes:
$r(\boldsymbol{z},\boldsymbol{q})=\mathrm{min}(\mathcal{R}(\boldsymbol{z},\boldsymbol{q}),\mathrm{clip}(\pi(\boldsymbol{z}\mid\boldsymbol{x},\boldsymbol{q}),1-\epsilon,1+\epsilon))$ (3)
where $\epsilon$ is a hyperparameter indicating the range of CLIP to be performed. After pruning, the extracted contextual content $g$ becomes the context $g^{{}^{\prime}}$, which is closely related to the reasoning objective $q$. We perform a pruning operation for each $g\in\mathcal{G}$ to obtain the pruned context space $\mathcal{G}^{{}^{\prime}}$.
3.3 Reasoning
After the re-organization process in Section 3.2, we get the re-organized context $g^{{}^{\prime}}\in\mathcal{G}^{{}^{\prime}}$. Then the re-organized context $g^{{}^{\prime}}$ can be used as a context alone or combined with the original context $c$ to get the final answer of $q$. The reasoning process can be defined as:
$o=f_{\theta}(g^{{}^{\prime}},\left[c\right],q,P_{r})$ (4)